# Well-Posed Geometric Boundary data in General Relativity, III: Conformal-volume data

### v1

@@ 2.4. Geometry at the corner
As noted in [2], the conformal factor $\varphi$ of the given boundary data $[g_{{\mathcal{C}}}]$ is determined at the corner $\Sigma$ by the initial data $g_{S}$. Consequently, the full metric $g_{\mu\nu}$ at $\Sigma$ is determined at the corner except possibly for the corner angle
@@
Let $\sigma$ denote a fixed background metric in the conformal class $[g_{{\mathcal{C}}}]$, so that $g_{{\mathcal{C}}}=\varphi^{2}\sigma$, where $[\sigma]$ is determined by the target data but $\varphi$ is not. As noted above, $\varphi|_{\Sigma}$ is determined by the initial and boundary data. Similarly, the initial velocity of $\varphi$ at $\Sigma$ is determined by the target data; in fact from [2], one has
@@ 3. Linear Analysis I
(3.1) $[g_{{\mathcal{C}}}]^{\prime}_{h}=\ring{h}=h^{\mathsf{T}}-\frac{\mathrm{tr}_{{\mathcal{C}}}h^{\mathsf{T}}}{n}g_{{\mathcal{C}}},$
@@
(3.6) $-\tfrac{n-1}{n}\Box_{\mathcal{C}}u-\tfrac{1}{n}R_{\mathcal{C}}u=Y({\mathcal{T}})+E(h),$
@@
There exists $\varepsilon_{0}>0$ such that, for any smooth metric $g$ on $\widetilde{U}$ which is $C^{\infty}$ $\varepsilon$-close, $\varepsilon\leq\varepsilon_{0}$, to a standard Minkowski corner metric $g_{\alpha_{0}}$ as in (2.5), the linearization $D\Phi_{g}^{H}$ at $g$ satisfies: given any $C^{\infty}$ target data $\tau^{\prime}\in T_{\tau}({\mathcal{T}}^{H})$ on $\widetilde{U}$, there exists a variation $h\in T_{g}Met(\widetilde{U})$, such that
@@
$-\tfrac{1}{2}\Box_{g}h^{\mathsf{T}}=F^{\mathsf{T}},$
@@
$-\tfrac{1}{2}\Box_{g}h(\nu)^{\mathsf{T}}=F(\nu)^{\mathsf{T}},$
@@
$-\tfrac{1}{2}\Box_{g}E_{1}^{\mathsf{T}}=-P^{\mathsf{T}}(h_{0}),$
@@
$-\tfrac{1}{2}\Box_{g}E_{1}(\nu)^{\mathsf{T}}=-P(\nu)^{\mathsf{T}}(h_{0})$
@@
(3.16) $\begin{split}-\tfrac{n-1}{n}\Box_{\mathcal{C}}\bar{u}-\tfrac{1}{n}R_{\mathcal{C}}\bar{u}=&R^{\prime}_{\ring{h}}-\mathrm{tr}(F-\delta^{*}\bar{V}^{\prime})+2(F-\delta^{*}\bar{V}^{\prime})(\nu,\nu)\\ &+2\langle A^{\prime}_{\bar{h}},A\rangle-2H_{\mathcal{C}}H^{\prime}_{\bar{h}}-2\langle\widehat{A}\circ A,\bar{h}\rangle+4\mathrm{Ric}(\nu,\nu^{\prime}_{\bar{h}})+\langle\mathrm{Ric},\bar{h}\rangle\\ &+\mathrm{tr}[(\delta^{*})^{\prime}_{\bar{h}}V]-2[(\delta^{*})^{\prime}_{\bar{h}}V](\nu,\nu),\\ \end{split}$
@@ 4. Energy estimates
(4.1) $\begin{split}||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}=[||F||_{H^{s}(\widetilde{U})}&+[||\ring{\gamma^{\prime}}||_{H^{s+1}({\mathcal{C}})}+||\mu^{\prime}||_{H^{s+1}({\mathcal{C}})}+||V^{\prime}_{{\mathcal{C}}}||_{H^{s}({\mathcal{C}})}]\\ &+[||\gamma^{\prime}_{S}||_{H^{s+1}(S_{0})}+||\kappa^{\prime}||_{H^{s}(S_{0})}+||\nu^{\prime}||_{H^{s}(S_{0})}+||V^{\prime}_{S}||_{H^{s}(S_{0})}].\end{split}$
@@
(4.3) $-\tfrac{1}{2}\Box_{g}h^{\mathsf{T}}+P^{\mathsf{T}}(h)=F^{\mathsf{T}},$
(4.4) $-\tfrac{1}{2}\Box_{g}h_{\nu\nu}+P_{\nu\nu}(h)=F_{\nu\nu},$
(4.5) $-\tfrac{1}{2}\Box_{g}h(\nu)^{\mathsf{T}}+P(\nu)^{\mathsf{T}}(h)=F(\nu)^{\mathsf{T}},$
@@
(4.6) $\begin{split}||u||_{\widetilde{H}^{s}(\Sigma_{t})}\leq C&[||F||_{H^{s-1}({\mathcal{C}}_{t})}+||\ring{h}||_{H^{s+1}({{\mathcal{C}}_{t})}}+||V^{\prime}_{h}||_{\bar{H}^{s}({\mathcal{C}}_{t})}+\varepsilon||h||_{\bar{H}^{s}({\mathcal{C}}_{t})}]\\ +&\ C[||u||_{H^{s}(\Sigma_{0})}+||\partial_{t}u||_{H^{s-1}(\Sigma_{0})}].\end{split}$
@@
$||u||_{H^{s}(\Sigma_{0})}+||\partial_{t}u||_{H^{s-1}(\Sigma_{0})}\leq C[||\gamma^{\prime}_{S}||_{H^{s+1/2}(S_{0})}+||\kappa^{\prime}||_{H^{s-1/2}(S_{0})}],$
@@
(4.7) $\begin{split}||V^{\prime}_{h}||_{\bar{H}^{s}({\mathcal{C}}_{t})}\leq C&[||F||_{H^{s}(\widetilde{U})}+||V^{\prime}_{{\mathcal{C}}}||_{H^{s}({\mathcal{C}}_{t})}+\varepsilon||h||_{H^{s}(\widetilde{U})}]\\ +&C[||V^{\prime}_{h}||_{H^{s}(S_{0})}+||\partial_{t}V^{\prime}_{h}||_{H^{s-1}(S_{0})}].\end{split}$
@@
$||\partial_{t}V^{\prime}_{h}||_{H^{s-1}(S_{0})}\leq C[||h_{S}||_{H^{s+1}(S_{0})}+||K^{\prime}_{h}||_{H^{s}(S_{0})}+||F||_{H^{s-1}(S_{0})}+||V^{\prime}_{h}||_{H^{s}(S_{0})}]\leq C||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}.$
@@
(4.8) $\begin{split}||h^{\mathsf{T}}||_{\bar{\mathcal{H}}^{s}(S_{t})}&\leq C[||h^{\mathsf{T}}||_{H^{s}({\mathcal{C}}_{t})}+\varepsilon||h||_{H^{s}(\widetilde{U}_{t})}+||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}]\\ &\leq C[\int_{0}^{t}[||\ring{h}||_{H^{s+1}({{\mathcal{C}}_{t^{\prime}})}}+||V^{\prime}_{{\mathcal{C}}}||_{H^{s}({\mathcal{C}}_{t^{\prime}})}.+\varepsilon||h||_{\bar{H}^{s}({\mathcal{C}}_{t^{\prime}})}]dt^{\prime}+\varepsilon||h||_{H^{s}(\widetilde{U}_{t})}+||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}],\\ \end{split}$
@@
(4.9) $\max_{t\in[0,1]}||h^{\mathsf{T}}||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq C[||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}+\varepsilon||h||_{\widehat{H}^{s}(\widetilde{U})}],$
where $||h||_{\widehat{H}^{s}(\widetilde{U})}:=||h||_{H^{s}(\widetilde{U})}+||h||_{\bar{H}^{s}({\mathcal{C}})}$.
@@
$\max_{t\in[0,1]}||h_{\nu\nu}||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq C[||h_{\nu\nu}||_{H^{s}({{\mathcal{C}})}}+\varepsilon||h||_{H^{s}(\widetilde{U})}+||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}],$
@@
(4.10) $\max_{t\in[0,1]}||h_{\nu\nu}||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq C[||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}+\varepsilon||h||_{\widehat{H}^{s}(\widetilde{U})}],$
@@
$||h(\nu)^{\mathsf{T}}||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq C[||\nabla_{\nu}h(\nu)^{\mathsf{T}}||_{H^{s-1/3}({{\mathcal{C}}_{t})}}+\varepsilon||h||_{H^{s}(\widetilde{U}_{t})}+||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}].$
@@
(4.11) $\begin{split}||h(\nu)^{\mathsf{T}}||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq&C[||\ring{h}||_{H^{s+2/3}({\mathcal{C}}_{t})}+||\mu^{\prime}_{h}||_{H^{s+2/3}({\mathcal{C}}_{t})}\\ &+||V^{\prime}_{{\mathcal{C}}}||_{H^{s-1/3}({\mathcal{C}})}+\varepsilon(||h||_{H^{s-1/3}({\mathcal{C}}_{t})}+||h||_{H^{s}(\widetilde{U}_{t})})+||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}].\end{split}$
Summing up the estimates (4.6)-(4.11), by choosing $\varepsilon$ sufficiently small depending on $g$, one may absorb the $\varepsilon||h||_{\bar{H}^{s}(\widetilde{U})}$ terms on the right side of these equations into the total sum on the left, using the fact that $|h||_{\bar{H}^{s}(\widetilde{U})}\leq\max_{t\in[0,1]}|h||_{\bar{\mathcal{H}}^{s}(S_{t})}$, giving then
(4.12) $\max_{t\in[0,1]}||h||_{\bar{\mathcal{H}}^{s}(S_{t})}\leq C||\tau^{\prime}||_{H^{s}({\mathcal{T}}^{H})}.$
@@ 5. Linear Analysis II
Then again $h_{0}$ has compact support in $\widetilde{U}_{0}$ for some time interval $[0,t_{0}]$ and as above extends to $M_{t_{0}}$. We relabel so that $t_{0}$ is a common time interval for all $h_{i}$, $i\geq 0$.
@@ 6. Nash-Moser theory
$||\widetilde{g}_{2}-\widetilde{g}_{1}||_{C^{\infty}}\leq\varepsilon,$
@@
We have shown in the proof of Theorem 1.1 above that $\mathbb{E}^{H}$ is a smooth, tame Frechet manifold. The group $\mathrm{Diff}_{0}(M)/\mathrm{Diff}_{1}(M)$ is a smooth tame Frechet Lie group (cf. [7, I.4.6]) which acts smoothly and freely on $\mathbb{E}^{H}$. It then follows from standard theory, cf. [7, Cor.III.2.5.2] that the quotient $\mathbb{E}^{H}/(\mathrm{Diff}_{0}(M)/\mathrm{Diff}_{1}(M))$ is a smooth, tame Frechet manifold. As noted in §1,
@@
indexed by $s$. The energy estimates in (5.2) hold on the $H^{s}$ completions of the domain and target spaces for suitable values of $s$, giving the existence of a tame inverse for the derivative $D\Phi^{H}$. One may then apply the finite derivative version of the Nash-Moser inverse function theorem due to Zehnder [13], or [14, Thm.6.3.3]; in the notation used there, the loss of derivative is $\gamma=3$. However, we will not carry out here the details that the hypotheses of [14, Thm.6.3.3] do hold.

### v2

@@ 2.4. Geometry at the corner
As noted in [2], the conformal factor $\varphi$ of the given boundary data $g_{{\mathcal{C}}}]$ is determined at the corner $\Sigma$ by the initial data $g_{S}$. Consequently, the full metric $g_{\mu\nu}$ at $\Sigma$ is determined at the corner except possibly for the corner angle
@@
Let $\sigma$ denote a fixed background metric in the conformal class $g_{{\mathcal{C}}}]$, so that $g_{{\mathcal{C}}}=\varphi^{2}\sigma$, where $\sigma]$ is determined by the target data but $\varphi$ is not. As noted above, $\varphi|_{\Sigma}$ is determined by the initial and boundary data. Similarly, the initial velocity of $\varphi$ at $\Sigma$ is determined by the target data; in fact from [2], one has
@@ 3. Linear Analysis I
(3.1) $g_{{\mathcal{C}}}]^{\prime}_{h}=\ring{h}=h^{\mathsf{T}}-\frac{\mathrm{tr}_{{\mathcal{C}}}h^{\mathsf{T}}}{n}g_{{\mathcal{C}}},$
@@
(3.6) $\tfrac{n-1}{n}\Box_{\mathcal{C}}u-\tfrac{1}{n}R_{\mathcal{C}}u=Y({\mathcal{T}})+E(h),$
@@
[…]
@@
$\tfrac{1}{2}\Box_{g}h^{\mathsf{T}}=F^{\mathsf{T}},$
@@
$\tfrac{1}{2}\Box_{g}h(\nu)^{\mathsf{T}}=F(\nu)^{\mathsf{T}},$
@@
$\tfrac{1}{2}\Box_{g}E_{1}^{\mathsf{T}}=-P^{\mathsf{T}}(h_{0}),$
@@
$\tfrac{1}{2}\Box_{g}E_{1}(\nu)^{\mathsf{T}}=-P(\nu)^{\mathsf{T}}(h_{0})$
@@
[…]
@@ 4. Energy estimates
[…]
@@
(4.3) $\tfrac{1}{2}\Box_{g}h^{\mathsf{T}}+P^{\mathsf{T}}(h)=F^{\mathsf{T}},$
(4.4) $\tfrac{1}{2}\Box_{g}h_{\nu\nu}+P_{\nu\nu}(h)=F_{\nu\nu},$
(4.5) $\tfrac{1}{2}\Box_{g}h(\nu)^{\mathsf{T}}+P(\nu)^{\mathsf{T}}(h)=F(\nu)^{\mathsf{T}},$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 5. Linear Analysis II
Then again $h_{0}$ has compact support in $\widetilde{U}_{0}$ for some time interval $0,t_{0}]$ and as above extends to $M_{t_{0}}$. We relabel so that $t_{0}$ is a common time interval for all $h_{i}$, $i\geq 0$.
@@ 6. Nash-Moser theory
$\widetilde{g}_{2}-\widetilde{g}_{1}||_{C^{\infty}}\leq\varepsilon,$
@@
We have shown in the proof of Theorem 1.1 above that $\mathbb{E}^{H}$ is a smooth, tame Frechet manifold. The group $\mathrm{Diff}_{0}(M)/\mathrm{Diff}_{1}(M)$ is a smooth tame Frechet Lie group (cf. [8, I.4.6]) which acts smoothly and freely on $\mathbb{E}^{H}$. It then follows from standard theory, cf. [8, Cor.III.2.5.2] that the quotient $\mathbb{E}^{H}/(\mathrm{Diff}_{0}(M)/\mathrm{Diff}_{1}(M))$ is a smooth, tame Frechet manifold. As noted in §1,
@@
indexed by $s$. The energy estimates in (5.2) hold on the $H^{s}$ completions of the domain and target spaces for suitable values of $s$, giving the existence of a tame inverse for the derivative $D\Phi^{H}$. One may then apply the finite derivative version of the Nash-Moser inverse function theorem due to Zehnder [15], or [16, Thm.6.3.3]; in the notation used there, the loss of derivative is $\gamma=3$. However, we will not carry out here the details that the hypotheses of [16, Thm.6.3.3] do hold.